Please help with the homework:
1. The scale of the map is 1cm:100m. Calculate the real distance if the map distance is 5,75 cm.

Answer by math_tutor2020(3817)


Answer: 575 m

Work Shown
1 cm : 100 m
5.75*1 cm : 5.75*100 m
5.75 cm : 575 m

Another approach
(1 cm)/(100 m) = (5.75 cm)/(x meters)
1/100 = 5.75/x
1*x = 100*5.75
x = 575
The result is a little over half a kilometer.